OpenStudy (amonoconnor):

Determine the Anti-Derivative of: "1/(3(x^(1/3)))" Any and all help is greatly appreciated! :)

OpenStudy (anonymous):

Can factor out the 1/3 and use power rule for integration on x^(-1/3).

OpenStudy (amonoconnor):

What did you get? This was from last night's assignment, but the answer key conflicts with what I got. I'm seeking a second option you could say :)

OpenStudy (anonymous):

\[\int x^{-1/3}dx = \frac{3}{2}x^{2/3} +c.\]

OpenStudy (anonymous):

a hard one(show me step to step solution) evaluate the definite integral by making a u substitution and integrating from u(a) to u(b) integral 0 to pi/2, cosx/(4+4sinx)^3dx